for the inverse demand function curve P=60-0.5Q. find the elasticity at p=10
1
Expert's answer
2017-12-19T08:57:06-0500
P = 60 - 0.5Q, Q = 120 - 2P, P = 10, Q = 120 - 2*10 = 100.
Ed = a*P/Q,
Ed = -2*10/100 = -0.2, so demand is inelastic.